OpenStudy (anonymous):

use the distributive property to multiply. 10(c+d)

OpenStudy (anonymous):

distributivity laws 1 & 2: A,B vectors a,b real numbers 1: (a+b)*A = a*A + b*A 2: a*(A+B) = a*A + a*B So if c,d are vectors, 10*(c+d) = 10*c + 10*d